A store owner want to develop a new snack mix by mixing chocolate and trail mix. How many pounds of chocolate costing $17.50 per pound should be mixed with trail mix costing $4.20 per pound to create a 31 mixture worth $8.92. (round to the nearest pound)
The owner needs to mix __________
pounds of chocolate and __________
pounds of trail mix.

Answer by addingup(3677)

17.50x + 4.20(31-x) = 8.92*31
17.50x + 130.20 - 4.20x = 276.52
13.30x = 146.32
x = 11 You need 11 chocolate and 31-11 = 20 trail mix